A typical jetliner lands at a speed of 71.5 m/s. If the plane travels at a constant speed of 71.5 m/s for 1.00 s, then apply the brakes and decelerates at a rate of -4.47 m/s2. What is the total displacement of the aircraft between touchdown and coming to rest?
1
Expert's answer
2020-10-23T11:52:36-0400

The distance covered when the aircraft was travelling with the speed "v_0 = 71.5m\/s" for "t_1 = 1s" is:


"d_1 = v_0t_1 = 71.5m/s\cdot 1s = 71.5m"

Time it took the aircraft to stop from "v_0 = 75.1m\/s" to "v = 0" under the acceleration "a = -4.47m\/s^2" is:


"t_2 = \dfrac{v_0}{|a|}"

The distance travelled in time "t_2" is:


"d_2 = v_0t_2 - \dfrac{|a|t_2^2}{2} = \dfrac{v_0^2}{|a|} - \dfrac{v_0^2}{2|a|} = \dfrac{v_0^2}{2|a|}"

Substituting numbers, obtain:


"d_2 = \dfrac{v_0^2}{2|a|} = \dfrac{71.5^2}{2\cdot 4.47} \approx 571.84m"

The total distance:


"d = d_1 + d_2 = 643.34m"

Answer. 643.34 m.